OpenStudy (anonymous):

Solve the following equation: a). y=7x + 4 b). y=6x + 14 y=9x - 6 2x - 9y =30

OpenStudy (anonymous):

Help!

OpenStudy (bobbyleary):

For a, since both functions are y = ..., this makes it easier to solve. You set both functions equal to each other. 7x + 4 = 9x - 6 Now solve for x. 10 = 2x x = 5 Now double check it by subsituting x = 5 each equation. y = 7(5) + 4 = 39 y = 9(5) - 6 = 39 The second one is one step harder, but give it a try. Post any issues you have and I will walk you through it.